OpenStudy (anonymous):

Solve the proportion: 4/(x+1) = 3/(x+2) there are two answers

OpenStudy (anonymous):

do you have any given answers to kinda help you with the questions& what math is this

OpenStudy (anonymous):

no i wish lol

OpenStudy (anonymous):

try gathering all x on the same side to start

hero (hero):

\[\frac{4}{x + 1} = \frac{3}{x + 2}\] Cross multiply \[4(x + 2) = 3(x + 1)\]

hero (hero):

Distribute: \[4x + 8 = 3x + 3\]

hero (hero):

Continue solving for x There can only be one correct value for x

OpenStudy (jdoe0001):

http://openstudy.com/study#/updates/52a22d0ee4b088e695c11550 <-- about the same thing

OpenStudy (anonymous):